OpenStudy (anonymous):

Express 9.39 · 10^9 seconds in terms of days?

OpenStudy (anonymous):

Can you show me the steps please?

OpenStudy (anonymous):

1 day = 24 hours = 24*60 minutes = 24*60*60 seconds

OpenStudy (anonymous):

does it help?

OpenStudy (mayankdevnani):

n 1 day there are 24 hours x 60 min/hour x 60 seconds/min = 86 400 seconds So divide 9.39 x 10^9 by 86 400 to determine the # of days there are = (9.39 x 10^9) ÷ (8.64 x 10^4) = (9.39 ÷ 8.64) x (10^9 ÷ 10^4) = 1.0868055556 x 10^5 = 108 681

OpenStudy (mayankdevnani):

ok @avieira

OpenStudy (mayankdevnani):

right @jiteshmeghwal9 and @sauravshakya

OpenStudy (jiteshmeghwal9):

yaa ! @mayankdevnani :)

OpenStudy (mayankdevnani):

thnx.. @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

yw :)

OpenStudy (anonymous):

Ohhhhh, okay!! I see what you have to do now, perfect. Thank you so much!!

OpenStudy (mayankdevnani):

welcome

OpenStudy (jiteshmeghwal9):

I think this question is based on reasoning :)

OpenStudy (jiteshmeghwal9):

or may be it is based on motion time etc.

OpenStudy (anonymous):